OpenStudy (annalee607):

The area of a regular Hexagon is 38cm (squared). What is the area of a regular hexagon with sides eight times as long? A)228cm(squared) B)304cm(squared) C)1,824cm(squared) D)2,432cm(squared)

OpenStudy (johnweldon1993):

Hint* Area of a hexagon is \(\large A = \frac{3\sqrt{3}}{2} s^2\) where 's' is a side length If that \(\large s\) now became \(\large 8s\) what would happen to the area?

OpenStudy (annalee607):

64

OpenStudy (johnweldon1993):

Right! It would increase by a factor of 64 So if the previous Area was \(\large 38cm^2\) then the new area would be \(\large 38 \times 64 = ? cm^2\)

OpenStudy (annalee607):

2,432?

OpenStudy (johnweldon1993):

\[\large \color \red{\checkmark}\]

OpenStudy (annalee607):

Thank you

OpenStudy (johnweldon1993):

Not a problem! Good job :)